45p+50n=27.75
70p+30n=36
How do I get my answer

Answer by scott8148(6628)

multiplying 1st eqn by -3/5 ___ -27p - 30n = -16.65

adding 2nd eqn ___ 43p = 19.35

dividing by 43 ___ p = .45

substituting ___ 70(.45) + 30n = 36.00 ___ 31.50 + 30n = 36.00 ___ 30n = 4.50 ___ n = .15